Shobo’s mother’s present age is six times Shobo’s present age. Shobo’s age five years from now will be one third of his mother’s present age. What are their present ages? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the present age of Shobo be x then age of her mother will be 6x.
Shobo's age after 5 years = x + 5
According to the question,
(x+5)=(1/3)×6x
⇒x+5=2x
⇒2x-x=5
⇒x=5
Present age of Shobo = x = 5 years
Present age of Shobo's mother = 6x = 30 years.

Thanks